# Null hypersurfaces in general relativity: Intrinsic symmetries and differential invariants

### v1

@@ Abstract
This paper investigates intrinsic Killing symmetries of null hypersurfaces $\mathcal{N}_{3}$ within the framework of general relativity. To this end we consider $\mathcal{N}_{3}$ as detached from the embedding spacetime and equipped with a degenerate metric of signature (0,+,+). As geometrical tools we use a triad calculus and differential invariants. Extending prior work, we present a classification of null hypersurfaces according to groups of motion up to the fourth order. For each type certain normal forms of the metric are given, and their invariants are listed. A discussion of horizons - defined as null hypersurfaces with vanishing shear and divergence - is included.
@@ 1 Introduction
Null hypersurfaces are embedded within the spacetime geometry, and geometrical quantities associated with this embedding - such as surface gravity and rigged null directions - play a crucial role in various applications. However, a clear geometric understanding can only be achieved by carefully separating intrinsic properties from embedding characteristics. In this spirit the null hypersurface can initially be considered as an independent object, detached from the ambient spacetime (see also [22]).
We adopt this strategy and investigate in this paper the Killing isometries of a three-dimensional standalone null hypersurface $\mathcal{N}_{3}$. A related study was carried out previously [13], but it was incomplete and contained inaccuracies. Here we correct and extend that analysis.
@@ 3 Differential invariants
$D^{\prime}\sigma^{\prime}=D\sigma+\sigma D\check{\lambda}-2i\check{\omega}D\sigma,$ (2.8m)
@@
$+i\delta\check{\omega}\Bigl(-\tau\frac{\partial f}{\partial\delta\tau}+(\bar{\tau}-\tau)\frac{\partial f}{\partial\delta\bar{\tau}}-\frac{\partial f}{\partial\bar{\tau}}-2\sigma\frac{\partial f}{\partial\delta\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}\Bigr)$
$+i\bar{\delta}\check{\omega}\Bigl(\bar{\tau}\frac{\partial f}{\partial\delta\tau}-\tau\frac{\partial f}{\partial\bar{\delta}\tau}+\bar{\tau}\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}+\frac{\partial f}{\partial\tau}+2\sigma\frac{\partial f}{\partial\bar{\delta\sigma}}+2\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}\Bigr)$
@@
$+\bar{\delta}\check{\kappa}\Bigl(-\rho\frac{\partial f}{\partial\bar{\delta}\tau}+\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}\Bigr)+\bar{\delta}\check{\bar{\kappa}}\Bigl(\sigma\frac{\partial f}{\partial\bar{\delta}\tau}-\rho\frac{\partial f}{\partial\delta\bar{\tau}}\Bigr)$
$-i\delta\delta\check{\omega}\frac{\partial f}{\partial\delta\bar{\tau}}+i\bar{\delta}\bar{\delta}\check{\omega}\frac{\partial f}{\partial\bar{\delta}\tau}+i\bar{\delta}\delta\check{\omega}\Bigl(\frac{\partial f}{\partial\delta\tau}-\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}\Bigr).$
@@
$\frac{\partial f}{\partial\bar{\delta}\tau}=0,~\frac{\partial f}{\partial\delta\bar{\tau}}=0,~\frac{\partial f}{\partial\delta\tau}=\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}.$ (2.8ae)
@@
II is itself an invariant of the system (written here with a non-vanishing rotation coefficient $\nu$). Obviously, we have to consider the following cases separately:
@@
$\zeta=2j-iI_{1}+I_{2},~\vartheta=2j-iI_{1}-I_{2}$ (2.8ay)
is a solution and represents an integral basis. In the case
$|\sigma|\neq 0,~\rho\neq 0,~I_{2}=0$
@@
itself is a new complex invariant for case $I_{2}=0$. As abbreviation we frequently use
@@
$+\bar{\delta}\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+\delta\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}+\bar{\delta}\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}+\sigma\frac{\partial f}{\partial\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}=0,$ (2.8bc)
$\sigma\frac{\partial f}{\partial D\sigma}+\bar{\sigma}\frac{\partial f}{\partial D\bar{\sigma}}=0,$ (2.8bd)
@@
$+3\delta\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}+\bar{\delta}\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}-2\sigma\frac{\partial f}{\partial\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}-\tau\frac{\partial f}{\partial\tau}+\bar{\tau}\frac{\partial f}{\partial\bar{\tau}}=0,$ (2.8be)
$-\frac{\partial f}{\partial\bar{\tau}}-2\sigma\frac{\partial f}{\partial\delta\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}=0,$ (2.8bf)
$-\frac{\partial f}{\partial\tau}+2\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}-2\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}=0,$ (2.8bg)
$\sigma\frac{\partial f}{\partial\delta\sigma}+\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}=0,$ (2.8bh)
$\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}=0,$ (2.8bi)
$D\sigma\frac{\partial f}{\partial\delta\sigma}+D\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}-\sigma\frac{\partial f}{\partial\tau}=0,$ (2.8bj)
$D\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+D\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}-\bar{\sigma}\frac{\partial f}{\partial\bar{\tau}}=0,$ (2.8bk)
$\frac{\partial f}{\partial\bar{\delta}\tau}=0,~\frac{\partial f}{\partial\delta\bar{\tau}}=0,~\frac{\partial f}{\partial\delta\tau}=\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}.$ (2.8bl)
@@
$\frac{\delta f}{\delta\bar{\delta}\bar{\sigma}}(4-I_{1}^{2})\frac{|\sigma|^{3}e^{-is}}{Ds}=0.$ (2.8bm)
@@
$N=e^{is/2}(\frac{\delta\sigma}{\sigma}-\frac{\delta\bar{\sigma}}{\bar{\sigma}}-4\bar{\tau})/(1+i)-ie^{-is/2}(\frac{\bar{\delta}\sigma}{\sigma}-\frac{\bar{\delta}\bar{\sigma}}{\bar{\sigma}}+4\tau)/(1+i).$ (2.8bn)
@@
$2\tau\frac{\partial f}{\partial\delta\tau}+\frac{\partial f}{\partial\bar{\tau}}=0,~\bar{\tau}(\frac{\partial f}{\partial\delta\tau}+\frac{\partial f}{\partial\bar{\delta}\bar{\tau}})+\frac{\partial f}{\partial\tau}=0,~-\tau\frac{\partial f}{\partial\tau}+\bar{\tau}\frac{\partial f}{\partial\bar{\tau}}=0,$~ (2.8bo)
$\frac{\partial f}{\partial\delta\bar{\tau}}=\frac{\partial f}{\partial\bar{\delta}\tau}=0,~\frac{\partial f}{\partial\delta\tau}=\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}.$ (2.8bp)
@@
Appendix $A$ provides a brief overview of the results of this section. One important point should be mentioned: There exist no unique set of invariants valid for all null hypersurfaces. Different classes of hypersurfaces have different sets of differential invariants.
@@ 6.4 Non-Abelian groups $G_{2}II$ with spacelike generators coplanar with the null direction
$j=\sqrt{1-1/k},~I_{1}=-2/\sqrt{k},~I_{2}=I_{3}=0,$ (2.8bl)
$~M=-i\frac{k_{,2}e^{is/2}}{(k-1)^{3/2}\sqrt{2k}},$ (2.8bm)
@@ 9 Final comments
$M=-2e^{is/2}\delta j/j,$ (2.8atauazbcbebhl)
$N=e^{is/2}(\frac{\delta\sigma}{\sigma}-\frac{\delta\bar{\sigma}}{\bar{\sigma}}-4\bar{\tau})/(1+i)-ie^{-is/2}(\frac{\bar{\delta}\sigma}{\sigma}-\frac{\bar{\delta}\bar{\sigma}}{\bar{\sigma}}+4\tau)/(1+i),$ (2.8atauazbcbebhm)
$K=\delta\tau+\bar{\delta}\bar{\tau}-2\tau\bar{\tau}.$ (2.8atauazbcbebhn)
We stress again: $j,I,J,K,M,N$ are invariants only, if certain conditions are satisfied for the rotation coefficients. These conditions can be taken from Table 1. E.g., $N$ is an invariant if and only if $|\sigma|\neq 0,~\rho=0$ and $|I_{1}|=2$.
@@
$E=a$ $X_{1}=\partial_{3}$ $j=\sqrt{1-c^{2}/(ab)}$ 2
$F=ce^{-x_{1}}$ $X_{2}=\partial_{2}$ $I_{1}=-2c/\sqrt{ab},~I_{2}=0$
$G=be^{-2x_{1}}$ $X_{3}=\partial_{1}+x_{3}\partial_{3}$ $I_{3}=0,~J=0,~M=0$
@@
$E=e^{-qx_{1}}\bigl(a_{1}+a_{2}\cos(px_{1})+a_{3}\sin(px_{1})\bigr)$ $X_{1}=\partial_{3}$ $j=$ 2
@@
$G=k(x_{2})e^{-2x_{1}}$ $X_{3}=2x_{3}\partial_{1}+x_{3}^{2}\partial_{3}$ $M=-ik_{,2}e^{is/2}/\sqrt{2k(k-1)^{3}}$
@@
$E=1$ $X_{1}=\partial_{3}$ $j=\frac{1}{2}$ 2
$F=e^{-x_{1}}$ $X_{2}=\partial_{1}+x_{3}\partial_{3}$ $I_{1}=-\frac{2}{\sqrt{k}},~I_{2}=0$
$G=ke^{-2x_{1}}$ $X_{3}=2x_{3}\partial_{1}+x_{3}^{2}\partial_{3}$ $J=M=0$

### v2

@@ Abstract
This paper investigates intrinsic Killing symmetries of null hypersurfaces $\mathcal{N}_{3}$ within the framework of general relativity. To this end we consider $\mathcal{N}_{3}$ as detached from the embedding spacetime and equipped with a degenerate metric of signature (0,+,+). As geometrical tools we use a triad calculus and differential invariants. Based on prior work, we present a classification of null hypersurfaces according to groups of motion up to the fourth order. For each type certain normal forms of the metric are given, and their invariants are listed. A discussion of horizons - defined as null hypersurfaces with vanishing shear and divergence - is included.
@@ 1 Introduction
Null hypersurfaces are embedded within the spacetime geometry, and geometrical quantities associated with this embedding - such as surface gravity and rigged null directions - play a crucial role in various applications. However, a clear geometric understanding can only be achieved by carefully separating intrinsic properties from embedding characteristics. Thus the null hypersurface can initially be considered as an independent object, detached from the ambient spacetime (see also [22]).
We adopt this strategy and investigate in this paper the Killing isometries of a three-dimensional standalone null hypersurface $\mathcal{N}_{3}$. A related study was carried out previously [13], but it contained many misprints and errors. Here we correct that analysis.
@@ 3 Differential invariants
$D^{\prime}\sigma^{\prime}=D\sigma+\sigma D\check{\lambda}-2i\check{\omega}D\sigma+2\check{\lambda}D\sigma,$ (2.8m)
@@
$+i\delta\check{\omega}\Bigl(-\tau\frac{\partial f}{\partial\delta\tau}+\bar{\tau}\frac{\partial f}{\partial\delta\bar{\tau}}-\frac{\partial f}{\partial\bar{\tau}}-2\sigma\frac{\partial f}{\partial\delta\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}\Bigr)$
$+i\bar{\delta}\check{\omega}\Bigl(-\tau\frac{\partial f}{\partial\bar{\delta}\tau}+\bar{\tau}\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}+\frac{\partial f}{\partial\tau}-2\sigma\frac{\partial f}{\partial\bar{\delta\sigma}}+2\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}\Bigr)$
@@
$+\bar{\delta}\check{\kappa}\Bigl(-\rho\frac{\partial f}{\partial\bar{\delta}\tau}+\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}\Bigr)+\bar{\delta}\check{\bar{\kappa}}\Bigl(\sigma\frac{\partial f}{\partial\bar{\delta}\tau}-\rho\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}\Bigr)$
$-i\delta\delta\check{\omega}\frac{\partial f}{\partial\delta\bar{\tau}}+i\bar{\delta}\bar{\delta}\check{\omega}\frac{\partial f}{\partial\bar{\delta}\tau}+i\delta\bar{\delta}\check{\omega}\frac{\partial f}{\partial\delta\tau}-i\bar{\delta}\delta\check{\omega}\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}.$
@@
$\frac{\partial f}{\partial\bar{\delta}\tau}=0,~\frac{\partial f}{\partial\delta\bar{\tau}}=0,~\frac{\partial f}{\partial\delta\tau}=0,~\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}=0.$ (2.8ae)
@@
II is itself an invariant of the system, written here with a non-vanishing rotation coefficient $\nu$. To simplify the calculations, we assumed $\nu=0$. In the invariants $J,L,M,N$ to be derived, $\nu$ only appears via $I$, thus no transformation to $\nu\neq 0$ is necessary.
Obviously, we have to consider the following cases separately:
@@
$\zeta=2j-iI_{1}+I_{2},~\vartheta=2j-iI_{1}-I_{2},~e^{is}=\sigma/|\sigma|$ (2.8ay)
is a solution and represents an integral basis.- In the case
$|\sigma|\neq 0,~\rho\neq 0,~I_{2}=0$,
@@
itself is a new complex invariant in the case $I_{2}=0$. As abbreviation we frequently use
@@
If also $I_{3}=0$ applies, there is a further invariant
$L=e^{is/2}(\frac{\delta\sigma}{\sigma}-\frac{\delta\rho}{\rho}-2\bar{\tau})+e^{-is/2}(j+i\sqrt{1-j^{2}})(\frac{\bar{\delta}\bar{\sigma}}{\bar{\sigma}}-\frac{\bar{\delta}\rho}{\rho}-2\tau).$ (2.8bc)
@@
$+\bar{\delta}\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+\delta\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}+\bar{\delta}\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}+\sigma\frac{\partial f}{\partial\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}=0,$ (2.8bd)
$\sigma\frac{\partial f}{\partial D\sigma}+\bar{\sigma}\frac{\partial f}{\partial D\bar{\sigma}}=0,$ (2.8be)
@@
$+3\delta\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}+\bar{\delta}\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}-2\sigma\frac{\partial f}{\partial\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}-\tau\frac{\partial f}{\partial\tau}+\bar{\tau}\frac{\partial f}{\partial\bar{\tau}}=0,$ (2.8bf)
$-\frac{\partial f}{\partial\bar{\tau}}-2\sigma\frac{\partial f}{\partial\delta\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}=0,$ (2.8bg)
$-\frac{\partial f}{\partial\tau}+2\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}-2\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}=0,$ (2.8bh)
$\sigma\frac{\partial f}{\partial\delta\sigma}+\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}=0,$ (2.8bi)
$\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}=0,$ (2.8bj)
$D\sigma\frac{\partial f}{\partial\delta\sigma}+D\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}-\sigma\frac{\partial f}{\partial\tau}=0,$ (2.8bk)
$D\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+D\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}-\bar{\sigma}\frac{\partial f}{\partial\bar{\tau}}=0,$ (2.8bl)
$\frac{\partial f}{\partial\bar{\delta}\tau}=0,~\frac{\partial f}{\partial\delta\bar{\tau}}=0,~\frac{\partial f}{\partial\delta\tau}=\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}.$ (2.8bm)
@@
$\frac{\delta f}{\delta\bar{\delta}\bar{\sigma}}(4-I_{1}^{2})\frac{|\sigma|^{3}e^{-is}}{Ds}=0.$ (2.8bn)
@@
$N=e^{is/2}(\frac{\delta\sigma}{\sigma}-\frac{\delta\bar{\sigma}}{\bar{\sigma}}-4\bar{\tau})/(1+i)-ie^{-is/2}(\frac{\bar{\delta}\sigma}{\sigma}-\frac{\bar{\delta}\bar{\sigma}}{\bar{\sigma}}+4\tau)/(1+i).$ (2.8bo)
@@
$2\tau\frac{\partial f}{\partial\delta\tau}+\frac{\partial f}{\partial\bar{\tau}}=0,~\bar{\tau}(\frac{\partial f}{\partial\delta\tau}+\frac{\partial f}{\partial\bar{\delta}\bar{\tau}})+\frac{\partial f}{\partial\tau}=0,~-\tau\frac{\partial f}{\partial\tau}+\bar{\tau}\frac{\partial f}{\partial\bar{\tau}}=0,$~ (2.8bp)
$\frac{\partial f}{\partial\delta\bar{\tau}}=\frac{\partial f}{\partial\bar{\delta}\tau}=0,~\frac{\partial f}{\partial\delta\tau}=\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}.$ (2.8bq)
@@
Appendix $A$ provides a brief overview of the results of this section. One important point should be mentioned: There exist no unique set of invariants valid for all null hypersurfaces. Different classes of hypersurfaces admit different sets of differential invariants.
@@ 6.4 Non-Abelian groups $G_{2}II$ with spacelike generators coplanar with the null direction
$j=\sqrt{1-1/k},~I_{1}=-2/\sqrt{k},~I_{2}=I_{3}=J=0,$ (2.8bl)
$L=\frac{k_{,2}e^{is/2}}{(k-1)\sqrt{2k}},~M=-i\frac{k_{,2}e^{is/2}}{(k-1)^{3/2}\sqrt{2k}},$ (2.8bm)
@@ 9 Final comments
$K=\delta\tau+\bar{\delta}\bar{\tau}-2\tau\bar{\tau},$ (2.8atauazbcbebhl)
$L=e^{is/2}\Bigl[\frac{\delta\sigma}{\sigma}-\frac{\delta\rho}{\rho}-2\bar{\tau}\Bigr]+e^{-is/2}(j+i\sqrt{1-j^{2}})\Bigl[\frac{\bar{\delta}\bar{\sigma}}{\bar{\sigma}}-\frac{\bar{\delta}\rho}{\rho}-2\tau\Bigr],$ (2.8atauazbcbebhm)
$M=-2e^{is/2}\delta j/j,$ (2.8atauazbcbebhn)
$N=e^{is/2}(\frac{\delta\sigma}{\sigma}-\frac{\delta\bar{\sigma}}{\bar{\sigma}}-4\bar{\tau})/(1+i)-ie^{-is/2}(\frac{\bar{\delta}\sigma}{\sigma}-\frac{\bar{\delta}\bar{\sigma}}{\bar{\sigma}}+4\tau)/(1+i).$ (2.8atauazbcbebho)
We stress again: $j,I,J,K,L,M,N$ are invariants only, if certain conditions are satisfied for the rotation coefficients. These conditions can be taken from Table 1. E.g., $N$ is an invariant if and only if $|\sigma|\neq 0,~\rho=0$ and $|I_{1}|=2$.
@@
$E=a$ $X_{1}=\partial_{3}$ $j=\sqrt{1-c^{2}/(ab)}$ 2b
$F=ce^{-x_{1}}$ $X_{2}=\partial_{2}$ $I_{1}=-2c/\sqrt{ab},~I_{2}=0,~I_{3}=0$
$G=be^{-2x_{1}}$ $X_{3}=\partial_{1}+x_{3}\partial_{3}$ J=0,L=0,M=0J=0,~L=0,~M=0
@@
$E=e^{-qx_{1}}\bigl(a_{1}+a_{2}\cos(px_{1})+a_{3}\sin(px_{1})\bigr)$ $X_{1}=\partial_{3}$ $j=$ 2a
@@
$G=k(x_{2})e^{-2x_{1}}$ $X_{3}=2x_{3}\partial_{1}+x_{3}^{2}\partial_{3}$ $L=k_{,2}e^{is/2}/\sqrt{2k(k-1)^{2}}$
$M=-ik_{,2}e^{is/2}/\sqrt{2k(k-1)^{3}}$
@@
$E=1$ $X_{1}=\partial_{3}$ $j=\sqrt{1-1/k}$ 2b
$F=e^{-x_{1}}$ $X_{2}=\partial_{1}+x_{3}\partial_{3}$ $I_{1}=-\frac{2}{\sqrt{k}},~I_{2}=0,~I_{3}=0$
$G=ke^{-2x_{1}}$ $X_{3}=2x_{3}\partial_{1}+x_{3}^{2}\partial_{3}$ $J=L=M=0$